How did they get 10.22 as the deviation? What did I do wrong?

Homework: Section 3.3 Save Score: 0 of 1 pt 5 of 9 (4 complete) HW Score: 44.44%, 4 of 9 pts 3.3.4-T Question Help The following data represent the number of people aged 25 to 64 years covered by health insurance (private or government) in 2003. Approximate the mean and standard deviation for age 25-34 20.2 35-44 33.4 45-54 32.8 55-64 23.5 Number (millions H44.92 years (Round to two decimal places as needed.) ơ: 20-21 years Round to two decimal places as needex Sorry, thats not correct. Sorry, your answer is not correct. Correct answer 10.22 Your answer 20.21 Similar Question Next Question Enter your answer in the answer box and then click Check Answer

My work:

L043 44 485.45 444983 0.8a2 0 11 d 니
Answer #1

The Table in excel

Age NumberF_{x} MidPoint(X) X*Fx F_{x}*X^{2}
25-34 20.2 29.5 595.9 17579.05
35-44 33.4 39.5 1319.3 52112.35
45-54 32.8 49.5 1623.6 80368.2
55-64 23.5 59.5 1398.25 83195.88
Sum= 109.9 Total\sum= 4937.05 233255.5
(\sum F_{x}*X)^{2}= 24374462.7

the mean is calculated as

\frac{\sum F_{x}*X}{n}

Mean calculated as 44.92

Standard Deviation is calculated as

\sigma = \sqrt{\frac{n{\sum (F_{x}*X^{2})}-{\sum (F_{x}*X)}^{2}}{n^{2}}}

and standard deviation as by putting all the data in formula

10.22